### v2

# Impact of the pre-equilibrium phase for the determination of nuclear geometry in high-energy isobar collisions

Abstract
Ultrarelativistic isobar collisions have been proposed as a useful tool to investigate nuclear structure. These systems are not created in equilibrium, rather undergo a prethermalization phase. In this phase, some of the initial structure information may be lost and additional effects introduced. The objective of this paper is to study this possibility in the extreme case of a “free-streaming” pre-equilibrium phase. We do this by computing estimators for ratios of various measured (or measurable) quantities (elliptic and triangular flows, mean transverse momentum and associated cumulants, correlators between elliptic or triangular flows and mean transverse momentum, symmetric cumulant and two-plane correlator) and study their sensitivity to the duration of the free-streaming phase. We find that the correlators between elliptic or triangular flows and mean transverse momentum, the so-called $\rho_{2}$ and $\rho_{3}$, are indeed sensitive to the duration of the free-streaming phase and that the normalized symmetric cumulant, $\epsilon NSC(2,3)$ might also depend on this duration.
pacs:
I Introduction
The successful application of the hydrodynamic framework to describe high energy heavy-ion collisions has opened a new door to study nuclear structure. Experimental data from the collaborations STAR Adamczyk et al. (2015); Abdallah et al. (2022); Xu (2022), ALICE Acharya et al. (2018, 2022), CMS Sirunyan et al. (2019) and ATLAS Aad et al. (2020, 2023) as well as theoretical studies Heinz and Kuhlman (2005); Shou et al. (2015); Xu et al. (2018); Li et al. (2020); Giacalone et al. (2021a, b); Jia (2022a, b); Jia et al. (2022); Zhang and Jia (2022); Jia and Zhang (2023); Xu et al. (2021a); Bally et al. (2022a); Nijs and van der Schee (2023); Zhao et al. (2023a), have indicated the importance of nuclear deformations and radial distributions of protons and neutrons in determining the measured multiplicity distribution and collective flow in these collisions.
Among deformed nuclei, the isobars ${}^{96}_{44}$u +${}^{96}_{44}$u and ${}^{96}_{40}$r +${}^{96}_{40}$r collisions at $\sqrt{s_{NN}}=200$ GeV studied by the STAR Collaboration Abdallah et al. (2022); Xu (2022) are of particular interest due to the similarities between the nuclei and the precision of the measurements. In these collisions, ${}^{96}_{44}$u exhibits higher particle production (by up to 10 percent), larger elliptic flow (by up to 3 percent), and smaller triangular flow (by 7 percent for central collisions) compared to ${}^{96}_{40}$r . These effects are relatively small, but still significant enough to be detected with high statistical precision and systematic uncertainties cancellation in ratios.
On the theory side, a large number of collision simulations is needed to achieve such high precision, and the process must be repeated when poorly understood nuclear geometry parameters are changed. Hydrodynamic-based simulations are computationally expensive, so one approach is to use the known mapping between initial conditions and anisotropic flow Gardim et al. (2012a, 2015a); Niemi et al. (2013, 2016); Fu (2015). Numerous studies have employed the aforementioned methodology, such as Jia (2022a, b); Jia et al. (2022); Zhang and Jia (2022). Nevertheless, a limited number of works have utilized hydrodynamic simulations Xu et al. (2021a); Nijs and van der Schee (2023); Zhao et al. (2023a).
State-of-the-art simulations of heavy-ion collisions are commonly based on a hybrid-chain of models, comprising four distinct stages: firstly, the generation of initial conditions; secondly, a prethermalization phase which serves to connect the initial conditions to the onset of hydrodynamic expansion; thirdly, the calculation of fluid expansion through the use of viscous relativistic hydrodynamics, and finally, a hadronic transport phase which occurs at the end of the hydrodynamic expansion. While the multiplicity distribution and collective flow are sensitive to the value of viscosity, as well as the inclusion of final state interactions and free streaming, it is expected that differences in the ratios of these observables only arise due to differences in nuclear structure. Although this assumption is reasonable, the precision of experimental data requires that it be tested. This has not be done systematically so far. An indirect study of viscosity and hadronic cascade effects using a transport model was recently conducted in Zhang et al. (2022).
A fundamental physical question in the preceding hydrodynamic description is the time after which the viscous hydrodynamic description becomes valid. Early investigations indicated that to reproduce the observed large hadron momentum anisotropy, this time must be smaller than about 1 fm Heinz and Kolb (2002). However, studies in the strong and weak coupling limits have not fully explained this rapid thermalization, which triggered new approaches to understand when hydrodynamics does apply (see e.g. the reviews Romatschke (2017); Florkowski et al. (2018); Romatschke and Romatschke (2019)). The objective of this paper is to investigate the effect of the prethermalization on the description of isobar collisions. This is done in the extreme scenario where immediately after production, partons are free streamed until the beginning of the hydrodynamic evolution. We note that such an approach has been used in various recent Bayesian analysis, where the most likely range of parameter values for the hydrodynamic scenario is sought Bernhard et al. (2019); Everett et al. (2021a, b); Parkkila et al. (2021); Nijs et al. (2021a, b); Liyanage et al. (2023). We perform our study using well known estimators for elliptic and triangular flows, but we also present estimators that have received little attention for isobars for quantities such as mean transverse momentum and associated cumulants, correlators between elliptic/triangular flow and mean transverse momentum, symmetric cumulant, mixed harmonic correlator.
II Setup of the simulations
Albeit many sophisticated models for nuclei exist from low-energy studies, the most popular way to model the nucleus in a ultra-relativistic collision is by sampling a Woods-Saxon distribution. The original distribution is spherically symmetric, hence not suitable for studying nuclei which are expected to have a deformed shape, such as ${}^{96}_{40}$r and ${}^{96}_{44}$u . The non-spherical shape is then taken into account by introducing an angular dependence in the nuclei radius, such as done in Jia (2022a); Giacalone et al. (2021b); Bally et al. (2022b). Hence, the probability $P(r,\,\theta,\,\varphi)$ of finding a nucleon in the position $\vec{r}=(r,\,\theta,\,\varphi)$ is proportional to the nuclear density profile
[…]
where the nuclear radius $R_{0}$, diffuseness $a$, quadrupole deformation parameter $\beta_{2}$, triaxiality deformation parameter $\gamma$ and octupole deformation parameter $\beta_{3}$ characterize the nuclei structure. It is expected that variations in these parameters are reflected in final-state observables. Here, motivated by the values for these parameters present in KIBÉDI and SPEAR (2002); Pritychenko et al. (2016); Zhang and Jia (2022) for ${}^{96}_{44}$u and ${}^{96}_{40}$r , we sample 6 distinct cases, where we progressively tune our parameters from a triaxial ${}^{96}_{44}$u nucleus (case 1) to the ${}^{96}_{40}$r nucleus (case 6). The parameters used are shown in Table 1
{tblr}
colspec=—Q[c,50pt]—Q[c,50pt]—Q[c,50pt]—Q[c,50pt]—Q[c,50pt]—Q[c,50pt]—, hlines = black, .5pt, vlines = black, .5pt Case Label & $\mathbf{R_{0}}$ (fm) $\mathbf{a}$ (fm) $\mathbf{\beta_{2}}$ $\mathbf{\beta_{3}}$ $\mathbf{\gamma}$
Case 1 5.09 0.46 0.16 0 $\pi/6$ [Uncaptioned image]
Case 2 5.09 0.46 0.16 0 0 [Uncaptioned image]
Case 3 5.09 0.46 0.16 0.20 0 [Uncaptioned image]
Case 4 5.09 0.46 0.06 0.20 0 [Uncaptioned image]
Case 5 5.09 0.52 0.06 0.20 0 [Uncaptioned image]
Case 6 5.02 0.52 0.06 0.20 0 [Uncaptioned image]
Table 1: Nuclear structure parameters employed for generating initial conditions of isobar collisions.
In Table 1, when going from one line to the next, only one parameter is changed. Cases 1 and 2 are similar except that the former is a triaxial ellipsoid and the later is a spheroid. An octupole deformation (pear shape) has been introduced when going from case 2 to 3. When going from case 3 to 4, the quadrupole deformation (ellipsoid shape) is reduced. Cases 4 and 5 are similar except that the later has a larger diffuseness. Lastly, the nuclear radius is reduced when going from case 5 to 6. These systematic modifications in the nuclear structure parameters allow to investigate eventual changes in the results when gradually moving from what is typically considered a Ruthenium nucleus (case 2) to what typically is used as a Zirconium nucleus (case 6). Effects related to “neutron-skin” have not been included in our simulations, therefore, protons and neutrons follow the same nuclear distribution. The reader may refer to Xu et al. (2021b) for a study on the neutron-skin effects.
Due to its flexibility and the ability to accept pre-generated (possibly deformed) nuclear configurations for projectile and target, the TRENTo model Moreland et al. (2015) has been used to generate initial conditions for our simulations. Two nuclear configurations (1) are randomly selected from a set of 50.000 pre-generated options and subsequently randomized in orientation. The Monte-Carlo Glauber model, where a value for the nucleon-nucleon cross-section $\sigma_{\rm NN}$ is needed, is then applied to identify which nucleons will participate in the collision, and to each participating nucleon, a density function $\rho_{i}(x,y,z)$ is assigned, namely a Gaussian function with standard deviation given by the nucleon width $w$. Each nucleon density is allowed to fluctuate according to a Gamma distribution with shape parameter $k$. By summing these densities and integrating over the $z$ direction for each nucleus, the participant thickness $T_{A,B}$ is obtained. These thicknesses are then used to calculate the energy density $\epsilon(x,y,\eta_{s}=0)$ as a reduced thickness between the two participant thicknesses given by
$\lim_{\tau\to 0^{+}}\tau\epsilon(x,y,\eta_{s}=0)=N\left[\frac{T_{A}(x,y)+T_{B}(x,y)}{2}\right]^{1/p}\,,$ (2)
where $N$ is a normalization constant and $p$ is called the reduced thickness parameter. Since in this work we will be dealing solely with ratios of quantities such as eccentricities and average radius of the system, the exact value of the normalization should not matter. However, in order to actually produce initial conditions with TRENTo , one needs to specify a value for its physical parameters ($N$, $w$, $\sigma_{\rm NN}$, $k$, $p$). In our simulations, these parameters were chosen to be the same as in Everett et al. (2021b) (because their analysis was done for both RHIC and LHC energies) and are outlined in Table 2 for completeness.
Parameter Symbol Value
Normalization (GeV/fm2) $N$ 5.73
Nucleon width (fm) $w$ 1.12
Cross-section (fm2) $\sigma_{\rm NN}$ 4.23
Fluctuation $k$ 0.907
Reduced thickness parameter $p$ 0.063
Table 2: TRENTo parameters used to generate initial conditionsEverett et al. (2021b).
After the collision takes place, one needs to specify how long the produced system will stay in the prethermalization phase. Given the model chosen in this work, such characteristic is controlled by the free-streaming time, $\tau_{FS}$. It is worth noting that at the moment all assumptions related to the modeling of pre-equilibrium phase are ad-hoc and such a quantity is currently chosen to have either a constant value or is allowed to carry a dependence with the collision centrality.
Regardless of the case, the value of $\tau_{FS}$ or its behavior with the collision centrality started to be determined (together with many other parameters needed in a complete hybrid hydrodynamic simulation) via Bayesian inference techniques Moreland et al. (2020); Bernhard et al. (2019); Everett et al. (2021b, a); Nijs et al. (2021a); Parkkila et al. (2021); Nijs and van der Schee (2022); Parkkila et al. (2022). This does not mean, however, that the main characteristics of this phase are well understood. In fact, it depends on the many particularities of each statistical analysis as different collaborations have obtained values for $\tau_{FS}$ that vary more than 50%: for instance, Parkkila et al. (2021) extracted $\tau_{FS}=0.901$ fm/c by considering Pb+Pb collisions at 5.02 TeV; on the other hand, Parkkila et al. (2022) reported $\tau_{FS}=0.71$ fm/c improving on the previous study by considering a simultaneous analysis of Pb+Pb data at 2.76 TeV and 5.02 TeV. When allowing the free-streaming to change with centrality as in Everett et al. (2021b), with parameters from the Grad viscous correction model, for Au+Au collisions at 200 GeV, $1.36\lesssim\tau_{FS}\,({\rm fm/c})\lesssim 1.51$, a variation of almost 50% when compared to the improved estimate from Parkkila et al. (2022). Lastly, a simultaneous analysis of p+Pb and Pb+Pb data at 5.02 TeV presented in Moreland et al. (2020) yielded $\tau_{FS}=0.37$ fm/c. While a smaller free-streaming time is expected for p+A collisions due to the system size after the collision takes place, this estimate (which is also applied to Pb+Pb collisions if one makes use of the model outlined in Moreland et al. (2020)) is $\sim 0.59\%$ smaller than the value quoted in Parkkila et al. (2021) which only considered Pb+Pb data at the same collision energy.
Given the current lack of a detailed understanding of the prethermalization phase and the large uncertainty in the extracted values for $\tau_{FS}$, in the present study we consider three different choices for the free-streaming time: (i) $\tau_{FS}=0.4$ fm/c, (ii) $\tau_{FS}=1.0$ fm/c and (iii) a centrality-dependent free-streaming time as implemented in Everett et al. (2021b): $\tau_{FS}=\tau_{R}(\{\epsilon\}/\{\epsilon_{R}\})^{\alpha}$, where $\{\epsilon\}$ denotes a (particular) average initial energy density in the transverse plane for a given collision, $\{\epsilon_{R}\}=4.0$ GeV/fm2, $\tau_{R}=1.46$ fm/c and $\alpha=0.031$ are constants that were adjusted by comparison with experimental data. These choices cover the range of values determined by recent Bayesian analysis and allow for a systematic investigation of eventual impacts that the prethermalization phase may have on quantities related to the system shape. In Fig. 1, the free-streaming time for the centrality-dependent case is shown as function of $\{\epsilon\}$ for Ru (very similar results hold for Zr). Centrality windows were estimated from ordering the total energy $E$ (integral on the transverse plane of the energy density in eq. (2) of 1 million TRENTo initial conditions in descending order for each colliding system. One sees that the free-streaming time varies between $\approx$ 1.35 and $\approx$ 1.49 fm in such a case.
Figure 1: Free-streaming time as function of $\{\epsilon\}$ in a mid-rapidity slice with $\tau_{R}=1.46$ fm/c, $\{\epsilon_{R}\}=4$ GeV/fm2 and $\alpha=0.031$. Centralities windows are also shown (see text). Very similar results are obtained for ${{}_{40}^{96}Zr}+{{}_{40}^{96}Zr}$.
III Interplay between free-streaming and nuclear geometry
Due to free-streaming, small scale features are expected to be smeared out and the uniform expansion results in decreased spatial asymmetries, i.e. smaller eccentricities $\varepsilon_{n}$, at the onset of the hydrodynamic phase, defined as
$\varepsilon_{n}$ =|$=\frac{|\{r^{n}e^{in\phi}\}|}{\{r^{n}\}}\,,$ (3)
[…]
$\varepsilon_{n}\{2\}$ $=\sqrt{\langle\varepsilon_{n}^{2}\rangle}\,,$ (4)
$\varepsilon_{n}\{4\}$ =$4]{2\langle\varepsilon_{n}^{2}\rangle^{2}-\langle\varepsilon_{n}^{4}\rangle}\,.$ (5)
Fig. 2 shows the same event at the initial time and at the end of the free-streaming. One sees that the energy density distribution becomes more rounded, with a larger average radius $R\equiv\sqrt{\{r^{2}\}}$ as time flows. Therefore, due to free-streaming, $R$ increases while the eccentricities decrease.
This is shown more precisely in Fig. 3. In this figure, the ratio between a given quantity computed after free-streaming and for the initial condition is used. It is a more effective mean of expressing the influence of the prethermalization phase. The following concise notation is used
${\rm ratio}[A]=\frac{\langle A\rangle_{FS}}{\langle A\rangle_{IC}}.$
Here and below, to determine centrality bin edges, all generated events are used, regardless of the (geometrical) case they belong to in Table 2. Such a procedure is different from the actual experiment, where centrality classification is based on the number of charged tracks together with information from Monte Carlo Glauber simulations for each collision system independently Abdallah et al. (2022). Results are in general shown in bins of width 1% in the 0-5% centrality window and then in bins of 5%; finer bins are used to $\rho_{2}$ and $\rho_{3}$, see below. When statistics demands, the 5 most central bins are merged in a single one of 0-5%. Centrality bin edges can be read from Table 3
Centrality Bin Energy Range (GeV) Centrality Bin Energy Range (GeV)
0-1% 700.00 - 466.10 10-15% 275.69 - 225.49
1-2% 466.10 - 444.98 15-20% 225.49 - 183.31
2-3% 444.98 - 426.64 20-25% 183.31 - 147.75
3-4% 426.64 - 409.73 25-30% 147.75 - 117.71
4-5% 409.73 - 336.08 30-35% 117.71 - 92.53
5-10% 336.08 - 275.69 35-40% 92.53 - 71.62
Table 3: Definition of the centrality bins in function of E.
In Fig. 3, curves end up grouped by how long the produced system free-streamed: for eccentricities, one group of curves corresponds to $\tau_{FS}=0.4$ fm/c, the middle one to $\tau_{FS}=1.0$ fm/c and the last one to the centrality-dependent $\tau_{FS}$, estimated to be approximately between 1.35 fm/c and 1.49 fm/c from Fig. 1. Free-streaming effects are increasingly stronger (meaning, larger deviations from 1) going from $R$, to $\epsilon_{2}$, then $\epsilon_{3}$ and finally $\epsilon_{4}$. Ratios for the 6 different geometrical configurations exhibit extremely similar shapes which mean they are affected fairly similarly by the free-streaming modeling of the prethermalization phase. A conformal expansion during the free-streaming is assumed. Since this model introduces no interactions as time goes by, in the case of a free-streaming velocity smaller than $c$ the system would have a slower expansion, reducing the effects of the prethermalization phase. An interacting model will introduce correlations on the system that will further slow down the eccentricity reductions. In this way, the magnitude of the effects shown here may be understood as an upper bound. For the subsequent hydrodynamic phase, reducing the free-streaming velocity is important for mitigating the artificial increase of the bulk pressure at the start of the expansion da Silva et al. (2023).
Figure 2: The mid-rapidity energy density slice of a ${}^{96}_{40}$r +${}^{96}_{40}$r event in the 20-30% centrality window is shown (left) initially and (right) after free-streaming.
We now concentrate on effects of varying the geometry, for a fixed free-streaming time. Results are shown for a given nuclear configuration (numbered from 1 to 5, with parameters from Table 1), relative to configuration 6 (expected to be more alike a Zr nucleus). For $\tau_{FS}=0.4$ fm/$c$, changes in the nuclear radius and eccentricities compared to their initial values are just a few percent (cf. Fig. 3) and ratios for a given configuration 1-5 with respect to 6 are essentially equal to the ones obtained at initial condition level, so not shown. Unless otherwise stated, we ran 10 million events minimum-bias to obtain the results shown.
Figure 3: Root mean square radius $R$ and eccentricities $\epsilon_{n}$ at the end of free-streaming phase divided by their initial values. Line styles correspond to different free-streaming times. One million minimum-bias events were used to generate this figure.
Figure 4: Left panels: Ratios between ellipticity computed for cases 1 through 5 and case 6, as function of total initial energy and with different free-streaming times. From top to bottom, the calculation is made for mean ellipticity, second and fourth order cumulants. Same line patterns as in Fig. 3 are used, with the addition of solid line to represent quantities computed for the initial conditions. Right panels: ratios between the curves of the left hand panel with free-streaming time and the ones for the initial conditions.
Figure 4 illustrates two effects: the first column represents the influence of distinct nuclear geometries, while the second column emphasizes the effect of free-streaming. The first column of Figure 4 shows ratios for the mean ellipticity employing the different nuclear configurations considered in this work. Configurations 1 and 2 have very similar ellipticity and differ only due to triaxiality $\gamma$ which should not affect this quantity. Configuration 3 shows a higher ellipticity than configuration 2 even though both configurations only differ by the inclusion of an octupole deformation $\beta_{3}$, indicating the non-trivial interplay between $\beta_{2}$ and $\beta_{3}$ discussed e.g. in Jia (2022a). The decrease in ellipticity from configuration 3 to 4 is expected due to the decrease of the quadrupole deformation. Configurations 4 and 5 have similar ellipticity, with ratios close to 1, indicating that this quantity is insensitive to nuclear radius and diffuseness. The differences in geometry are more apparent for the most central collisions.
For the second order cumulant ratios: $\epsilon_{2}\{2\}_{n}/\epsilon_{2}\{2\}_{6}=\langle\epsilon_{2}\rangle_{n}/\langle\epsilon_{2}\rangle_{6}\times\sqrt{1+(\sigma^{2}_{n}/\langle\epsilon_{2}\rangle^{2}_{n})}^{2}\\ /\sqrt{1+(\sigma^{2}_{6}/\langle\epsilon_{2}\rangle^{2}_{6})^{2}}$. We have checked that the quantities $\sigma^{2}_{n}/\langle\epsilon_{2}\rangle^{2}_{n}$ are nearly equal for n=1-6, so it explains the similarity between the upper and lower panels. The ratios of $\epsilon_{2}\{4\}$ are also shown, even though the error bars are large, because it has been argued that the comparison of $v_{2}\{2\}$ and $v_{2}\{4\}$ or $\epsilon_{2}\{2\}$ and $\epsilon_{2}\{4\}$ for isobars might yield information on nuclear structure Jia et al. (2023). The inset illustrates the effect of changing the diffuseness $a$.
[…]
Figure 5 shows the triangularity as a function of the total initial energy $E$ for different choices of the free-streaming times. On the left, by comparing the different nuclear configurations it is possible to verify that this quantity is mainly affected by the octupole deformation $\beta_{3}$ (as expected) and the diffuseness. The effect is more prominent in central collisions, where the nucleus overlap more and their nuclear geometry becomes increasingly important. Other details such as the nuclear radius, triaxiality, and quadrupole deformation do not affect this quantity. From the right hand side, it can be seen that free-streaming has larger effects for triangularity than for ellipticity for cases 1 and 2 (divided by that of case 6): there is a larger splitting for the two different free-streaming times, almost 1% for most centralities, which is comparable with the error bars in experimental triangular flow ratios. For cases 3-5, the effect is a little less pronounced. In any case, the modeling and duration of the prethermalization phase may play a role in a very precise determination of the octopole deformation.
Figure 5: Ratios of triangularities, similar to Fig. 4.
Figure 6 shows a higher-order eccentricity, $\varepsilon_{4}$, as function of the total initial energy $E$ for different free-streaming times. When going from case 4 to case 5, for central collisions, the dependence on the diffuseness parameter is 3% for central collisions and larger for non central ones. In line with results obtained for triangularity, a progressively larger splitting between the results with different free-streaming times is observed.
Figure 6: Ratios of higher order eccentricities, similar to Fig. 4.
[…]
Figure 7: Ratios of root mean square of radius, in a similar fashion as done in Fig. 4 for ellipticity.
Given that a good mapping is expected between not just $v_{2}$, $v_{3}$ and $\epsilon_{2}$, $\epsilon_{3}$, but also between the event plane angles and initial state symmetry plane angles $\Phi_{2}$, $\Phi_{3}$, it can be used to study the normalized symmetric cumulant NSC(2,3) Gardim et al. (2017),
$\varepsilon NSC(2,3)=\frac{\langle\varepsilon_{2}^{2}\varepsilon_{3}^{2}\rangle-\langle\varepsilon_{2}^{2}\rangle\langle\varepsilon_{3}^{2}\rangle}{\langle\varepsilon_{2}^{2}\rangle\langle\varepsilon_{3}^{2}\rangle}\,$ (6)
as well as the two-plane correlator
$\langle\cos 6(\Phi_{2}-\Phi_{3})\rangle=\frac{\langle\epsilon_{2}^{3}\epsilon_{3}^{2}\cos 6(\Phi_{2}-\Phi_{3})\rangle}{\sqrt{\langle\epsilon_{2}^{6}\rangle\langle\epsilon_{3}^{4}\rangle}}\,.$ (7)
Figure 8: Normalized symmetric cumulant $\epsilon NSC(2,3)$ computed from initial conditions for the 6 geometrical cases.
In Fig. (8), for the normalized symmetric cumulant computed directly from the initial conditions, cases 1, 2, 3 are similar and 4 and 5 are fairly equal. Free streaming seems to have a strong effect, and might change once or twice the sign of the ratio, however error bars are large, so more statistics would be necessary to confirm this. The changes of sign can be traced back to slightly different free-streaming time dependencies for each nuclear geometry considered: as the system evolves in time, the individual quantities $\varepsilon NSC(2,3)$ appearing in this ratio becomes less negative, with each of them crossing zero at different energy values. Case 6 only shows a single change of sign (positive to negative) that always happens for values of energy lower than the ones shown in our plots, meanwhile, cases 2 and 3 for instance, present two clear changes of sign (positive to negative and back to positive), with the second change happening in the most central bin considered. The ratio of two-plane correlators, fluctuates around zero and suffer from lack of statistics so is not shown. Another interesting angle correlator for which data have been obtained is $ac2(3)=\langle v_{2}^{2}v_{4}\cos 4(\Psi_{2}-\Psi_{4})\rangle$. Since it cannot be predicted for all centralities with just $\varepsilon_{4}$ Gardim et al. (2012b, 2015b), we leave it to a future study from hydrodynamics. We note however that it has a good potential to discriminate geometry in central collisions Zhao et al. (2023b).
The observables presented so far, except perhaps $\varepsilon NSC(2,3)$, exhibit minimal sensitivity to the effects of free-streaming and triaxiality. One observable known to be sensitive to triaxiality is the Pearson correlator between event anisotropic flow and event mean transverse momentum $[p_{T}]$ in a given centrality bin Bally et al. (2022a); Aad et al. (2023). This correlator has been identified as a strong candidate for assessing the effects on triaxility in isobars Jia (2022b). Previous studies were limited to $n=2$ and utilized estimators that lack proper correlation with mean transverse momentum: Ref. Giacalone et al. (2021c) showed that a good proxy to estimate this type of correlator is to replace the mean momentum by the ratio between total energy and total entropy in a very narrow centrality bin and both $v_{2}\propto\varepsilon_{2}$ and $v_{3}\propto\varepsilon_{3}$ can be considered. Therefore we calculate the following Pearson correlator:
$\rho_{n}=\frac{\langle v_{n}^{2}[p_{T}]\rangle-\langle v_{n}^{2}\rangle\langle[p_{T}]\rangle}{\sigma_{v_{n}^{2}}\sigma_{[p_{T}]}}\approx\frac{\langle\varepsilon_{n}^{2}E/S\rangle-\langle\varepsilon_{n}^{2}\rangle\langle E/S\rangle}{\sigma_{\varepsilon_{n}^{2}}\sigma_{E/S}}\,,$ (8)
where the averages are taken in centrality bins of 0.25% centrality and $\sigma_{\varepsilon_{n}^{2}}$ and $\sigma_{E/S}$ are the standard deviation of $\varepsilon_{n}$ and $E/S$ in those centrality bins, respectively.
[…]
Our results confirm that the effect the triaxiality parameter $\gamma$ is particularly clear when comparing configurations 1 and 2 for $\rho_{2}$ in the top panel of Fig. 9. We also report that the splitting between the results of these two cases increases with the free-streaming time, starting from $\sim 0.3$ at the initial condition level and reaching $\sim 0.5$ when considering a centrality dependent free-streaming duration. Comparing cases 2 and 3, shows that the value of $\beta_{3}$ has a strong effect, lowering the correlation in the $\sim 5\%-13\%$ centrality range. Moreover, one verifies that the free-streaming time plays an interesting role, inducing a single or a double sign change depending on how long the system stays in the prethermalization phase. While we still lack final experimental results for this correlator, such a signature may eventually help to constrain the duration of the pre-hydrodynamic phase in isobar collisions. Next, reducing the quadrupole deformation while keeping the octupole deformation fixed shifts the correlation to values around unity with a much reduced free-streaming time dependence. For the case 5, the result is brought even closer to unity. Using these results together with the ones presented in Fig. 4, 5 indicates that it might be possible to constrain $\beta_{2}$, $\beta_{3}$ and $\gamma$ simultaneously from observations of anisotropic flow and the correlation of elliptical flow with mean transverse momentum.
We have checked that using different estimators for $[p_{T}]$, the effect of the free-streaming time is still substantial.
Moving to $\rho_{3}$, and assuming that the estimator (8) is still a good proxy for the final state observable (as is the case for high-energy Pb+Pb collisions Giacalone et al. (2021c)), one sees that while the dependence on the triaxiality parameter becomes weaker, the effect of free-streaming time is still visible for cases 1-3. In addition, adding an octupole deformation on top of a quadrupole deformation (case 3) shifts the correlation to even larger values in central collisions, providing a clear separation with cases 1 and 2 where only the quadrupole deformation is present. Cases 4, 5 and 6 all have similar $\rho_{3}$, indicating that this observable is independent of diffuseness and distribution radius in the centrality range considered.
Figure 9: Top panel: Pearson correlation between eccentricity squared and the ratio of total energy by total entropy as function of total initial energy for initial conditions. Lower panel: The same, but replacing ellipticity by triangularity in the $\rho_{n}$ correlator.
Figure 10: Expected value for the distribution of E/S (in a centrality bin) for case n divided by that of case 6 as function of centrality.
Another possibility to determine the geometry is the study of the distribution of event mean transverse momentum in a given centrality bin, $[p_{T}]$. As argued above, a good estimator for $[p_{T}]$ is E/S. It is expected that a good estimator for $\langle\delta[p_{T}]^{p}\rangle/\langle[p_{T}]\rangle^{p}$ would be $\langle(\delta(E/S))^{p}\rangle/\langle E/S\rangle^{p}$ where $\delta O\equiv O-\langle O\rangle$ (as far as we know this has not been tested throughout). The ratio $\langle E/S\rangle_{n}/\langle E/S\rangle_{6}$ as function of centrality is shown in Fig. 10. It is computed directly from the initial conditions (as E/S is conserved) and exhibit a clear ordering: cases n=1-2 have the highest values for the ratio, above 1, and do not depend on the triaxiality. Case 3 is also above 1 but introducing $\beta_{3}$ has decreased the ratio. For case 4, decreasing $\beta_{2}$ increases the ratio. For case 5, increasing $a$ decreases the ratio. However, all these effects are quite small. Fig. 11 is more interesting: the ratios $\left(\langle(\delta(E/S))^{2}\rangle/\langle E/S\rangle^{2}\right)_{n}/\left(\langle(\delta(E/S))^{2}\rangle/\langle E/S\rangle^{2}\right)_{6}$ are shown and larger effects are observed. Cases n=1-2 have the most extreme values for the ratio, below 0.98 and above 1.01, and do not depend on the triaxiality. Case 3 also reaches above 1.01 but introducing $\beta_{3}$ has pushed the ratio towards 1 in general (by about 2%). For case 4, decreasing $\beta_{2}$ decreases the ratio. For case 5, increasing $a$ pushes the ratio very near 1 except for very central collisions. This quantity might be useful to further constrain the value of $\beta_{2}$ given that curves with small values of this parameter are well separated from those with higher values for central collisions. Finally in Fig. 12, $\left(\langle(\delta(E/S))^{3}\rangle/\langle E/S\rangle^{3}\right)_{n}/\left(\langle(\delta(E/S))^{3}\rangle/\langle E/S\rangle^{3}\right)_{6}$ is shown. It has been claimed Jia (2022b) that this quantity could also be used to constrain $\gamma$, but the effect seems very small. Though the error bars are big, some geometrical effects can be seen, for example, introducing $\beta_{3}$ increases the ratio and diminishing $\beta_{2}$ decreases it.
In Jia (2022a), some quantities comparable to those we computed have be obtained as well but with a different estimator for $[p_{T}]$. As a general rule, the effect of geometry is qualitatively similar.
Figure 11: Same as Fig. 10 for the variance.
Figure 12: Same as Fig. 10 for the third cumulant.
IV Conclusion
We investigated the effects of a prethermalization phase in isobar collision systems (namely ${}^{96}_{44}$u +${}^{96}_{44}$u and ${}^{96}_{40}$r +${}^{96}_{40}$r ), using free-streaming as an extreme case. The main effects anticipated were observed: the smearing out of small-scale features and reduction of spatial anisotropies (which get transferred to momentum distribution anisotropy). Despite the absolute magnitude of these effects being large, it is not the same when doing ratios for a given $\tau_{FS}$. In this case, for ratios of ellipticity and rms radii, differences are less than 1% and unlikely to be relevant when comparing to data. Ratios of higher order eccentricities, however, show differences that can be significant when comparing with precision measurements such as the ones recorded for isobar collisions. Within our statistics, not much can be learned from the two-plane correlator for elliptic and triangular flows but interestingly the normalized symmetric cumulant seems sensitive to the duration of free-streaming. These studies of eccentricities yield information on the parameters $\beta_{2}$, $\beta_{3}$ and $a$.
Furthermore, we studied the effects that nuclear geometry could have on the correlation between anisotropic flow and mean momentum to get information on $\gamma$. In Giacalone et al. (2021c), the estimator $\rho(\epsilon_{2}^{2},E/S)$ was shown to better reproduce the hydrodynamic correlator $\rho(v_{2}^{2},[p_{T}])$. Here for the first time, it is used to study isobar collisions. A several percent difference between the cases n=1 and 2 (with and without triaxiality) is indeed observed for $\rho_{2}$. In addition, and this is the main result as far as free-streaming effects are concerned, the duration of the prethermalization phase, regulated by $\tau_{FS}$, has a noticeable and diverse effect on these ratios of correlators depending on the description of the nuclear geometry: for $\rho_{2}$, it is able to enhance the separation between results with and without triaxiality effects (cases 1 and 2); for $\rho_{3}$ it generates a non-zero correlation between $\varepsilon_{3}$ and $E/S$ (proxies for $v_{3}$ and the mean transverse momentum, respectively) for the same cases. Another distinct effect for $\rho_{2}$ appears for case 3 (where in addition to the quadrupole deformation of cases 1 and 2, an octupole deformation is introduced), a double sign change for the largest free-streaming times ($1.35\lesssim\tau_{FS}\,({\rm fm/c})\lesssim 1.49$) appears, which can be contrasted to no or perhaps one sign change for initial condition and free-streaming time equal to $1$ fm/c. If verified experimentally, these features provide another tests of the predictive power of state-of-the-art hybrid hydrodynamic models Giannini et al. (2023).
The analysis with $E/S$ as estimator for $[p_{T}]$ is also used to study the distribution of $[p_{T}]$ and its first momenta, for initial conditions only (because $E/S$ is conserved during free-streaming) for the first time. The ratio of variances are most promising to constrain $\beta_{2}$ in addition to the eccentricities analysis. The ratio of third order cumulant which in principle could yield information on $\gamma$ as well, does not seem to lead to a strong effect.
In a future work, we will test how good the estimators presented here are, by comparing with actual hydrodynamic simulations. In turn, this will permit to select the best estimators and use them to perform a bayesian analysis to extract nuclear geometry.
V Acknowledgements
F.G.G. was supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq grant 306762/2021-8). K.P.P. and W.M.S. acknowledge support from Fundação de Amparo à Pesquisa do Estado de São Paulo (respectively via grants 2020/15893-4 and 2021/01670-6, 2022/11842-1). All authors acknowledge support from Fundação de Amparo à Pesquisa do Estado de São Paulo (FAPESP grant 2018/24720-6) and project INCT-FNA Proc. No. 464898/2014-5. A.V.G. has been partially supported by CNPq. The authors thank the HPC resources provided by Information Technology Superintendence (HPC-STI) of University of São Paulo (Aguia cluster). The authors acknowledge the National Laboratory for Scientific Computing (LNCC/MCTI, Brazil), through project COL-ISO as well as the ambassador program (UFGD), subproject FCNAE, which providing HPC resources from the SDumont supercomputer URL: http://sdumont.lncc.br. The authors thank the coordinators of the EMMI task force ”Nuclear physics confronts relativistic collisions of isobars”, G.Giacalone, J. Jia, V. Somà, Y. Zhou, who first suggested this line of research.

### v4

@@
-Impact of the pre-equilibrium phase for the determination of nuclear geometry in high-energy isobar collisions
+Impact of the pre-equilibrium stage for the determination of nuclear geometry in high-energy isobar collisions
 Abstract
 Ultrarelativistic isobar collisions have been proposed as a useful tool to investigate nuclear structure.
-These systems are not created in equilibrium, rather undergo a prethermalization phase.
-In this phase, some of the initial structure information may be lost and additional effects introduced.
-The objective of this paper is to study this possibility in the extreme case of a “free-streaming” pre-equilibrium phase.
-We do this by computing estimators for ratios of various measured (or measurable) quantities (elliptic and triangular flows, mean transverse momentum and associated cumulants, correlators between elliptic or triangular flows and mean transverse momentum, symmetric cumulant and two-plane correlator) and study their sensitivity to the duration of the free-streaming phase.
-We find that the correlators between elliptic or triangular flows and mean transverse momentum, the so-called $\rho_{2}$ and $\rho_{3}$, are indeed sensitive to the duration of the free-streaming phase and that the normalized symmetric cumulant, $\epsilon NSC(2,3)$ might also depend on this duration.
+These systems are not created in equilibrium, rather undergo a pre-thermalization stage.
+In this stage, some of the initial structure information may be lost and additional effects introduced.
+The objective of this paper is to study this possibility in the extreme case of a “free-streaming” pre-equilibrium stage.
+We do this by computing estimators for ratios of various measured (or measurable) quantities (elliptic and triangular flows, mean transverse momentum and associated cumulants, correlators between elliptic or triangular flows and mean transverse momentum, symmetric cumulant and two-plane correlator) and study their sensitivity to the duration of the free-streaming stage.
+We find that the correlators between elliptic or triangular flows and mean transverse momentum, the so-called $\rho_{2}$ and $\rho_{3}$, are indeed sensitive to the duration of the free-streaming stage and that the normalized symmetric cumulant, $\varepsilon NSC(2,3)$ might also depend on this duration.
 pacs:
@@
 The successful application of the hydrodynamic framework to describe high energy heavy-ion collisions has opened a new door to study nuclear structure.
-Experimental data from the collaborations STAR Adamczyk et al.
-(2015); Abdallah et al.
-(2022); Xu (2022), ALICE Acharya et al.
-(2018, 2022), CMS Sirunyan et al.
-(2019) and ATLAS Aad et al.
-(2020, 2023) as well as theoretical studies Heinz and Kuhlman (2005); Shou et al.
-(2015); Xu et al.
-(2018); Li et al.
-(2020); Giacalone et al.
-(2021a, b); Jia (2022a, b); Jia et al.
-(2022); Zhang and Jia (2022); Jia and Zhang (2023); Xu et al.
-(2021a); Bally et al.
-(2022a); Nijs and van der Schee (2023); Zhao et al.
-(2023a), have indicated the importance of nuclear deformations and radial distributions of protons and neutrons in determining the measured multiplicity distribution and collective flow in these collisions.
-Among deformed nuclei, the isobars ${}^{96}_{44}$u +${}^{96}_{44}$u and ${}^{96}_{40}$r +${}^{96}_{40}$r collisions at $\sqrt{s_{NN}}=200$ GeV studied by the STAR Collaboration Abdallah et al.
-(2022); Xu (2022) are of particular interest due to the similarities between the nuclei and the precision of the measurements.
+Experimental data from the collaborations STAR [1, 2, 3], ALICE [4, 5], CMS [6] and ATLAS [7, 8] as well as theoretical studies [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23], have indicated the importance of nuclear deformations and radial distributions of protons and neutrons in determining the measured multiplicity distribution and collective flow in these collisions.
+Among deformed nuclei, the isobars ${}^{96}_{44}$u +${}^{96}_{44}$u and ${}^{96}_{40}$r +${}^{96}_{40}$r collisions at $\sqrt{s_{NN}}=200$ GeV studied by the STAR Collaboration [2, 3] are of particular interest due to the similarities between the nuclei and the precision of the measurements.
 In these collisions, ${}^{96}_{44}$u exhibits higher particle production (by up to 10 percent), larger elliptic flow (by up to 3 percent), and smaller triangular flow (by 7 percent for central collisions) compared to ${}^{96}_{40}$r .
@@
 On the theory side, a large number of collision simulations is needed to achieve such high precision, and the process must be repeated when poorly understood nuclear geometry parameters are changed.
-Hydrodynamic-based simulations are computationally expensive, so one approach is to use the known mapping between initial conditions and anisotropic flow Gardim et al.
-(2012a, 2015a); Niemi et al.
-(2013, 2016); Fu (2015).
-Numerous studies have employed the aforementioned methodology, such as Jia (2022a, b); Jia et al.
-(2022); Zhang and Jia (2022).
-Nevertheless, a limited number of works have utilized hydrodynamic simulations Xu et al.
-(2021a); Nijs and van der Schee (2023); Zhao et al.
-(2023a).
-State-of-the-art simulations of heavy-ion collisions are commonly based on a hybrid-chain of models, comprising four distinct stages: firstly, the generation of initial conditions; secondly, a prethermalization phase which serves to connect the initial conditions to the onset of hydrodynamic expansion; thirdly, the calculation of fluid expansion through the use of viscous relativistic hydrodynamics, and finally, a hadronic transport phase which occurs at the end of the hydrodynamic expansion.
+Hydrodynamic-based simulations are computationally expensive, so one approach is to use the known mapping between initial conditions and anisotropic flow [47, 24, 48, 25, 26].
+Numerous studies have employed the aforementioned methodology, such as [15, 16, 17, 18].
+Nevertheless, a limited number of works have utilized hydrodynamic simulations [20, 22, 23].
+State-of-the-art simulations of heavy-ion collisions are commonly based on a hybrid-chain of models, comprising four distinct stages: firstly, the generation of initial conditions; secondly, a pre-thermalization stage which serves to connect the initial conditions to the onset of hydrodynamic expansion; thirdly, the calculation of fluid expansion through the use of viscous relativistic hydrodynamics, and finally, a hadronic transport phase which occurs at the end of the hydrodynamic expansion.
 While the multiplicity distribution and collective flow are sensitive to the value of viscosity, as well as the inclusion of final state interactions and free streaming, it is expected that differences in the ratios of these observables only arise due to differences in nuclear structure.
@@
 This has not be done systematically so far.
-An indirect study of viscosity and hadronic cascade effects using a transport model was recently conducted in Zhang et al.
-(2022).
+An indirect study of viscosity and hadronic cascade effects using a transport model was recently conducted in [27].
 A fundamental physical question in the preceding hydrodynamic description is the time after which the viscous hydrodynamic description becomes valid.
-Early investigations indicated that to reproduce the observed large hadron momentum anisotropy, this time must be smaller than about 1 fm Heinz and Kolb (2002).
+Early investigations indicated that to reproduce the observed large hadron momentum anisotropy, this time must be smaller than about 1 fm [28].
 However, studies in the strong and weak coupling limits have not fully explained this rapid thermalization, which triggered new approaches to understand when hydrodynamics does apply (see e.g.
-the reviews Romatschke (2017); Florkowski et al.
-(2018); Romatschke and Romatschke (2019)).
-The objective of this paper is to investigate the effect of the prethermalization on the description of isobar collisions.
+the reviews [29, 30, 31]).
+The objective of this paper is to investigate the effect of the pre-thermalization on the description of isobar collisions.
 This is done in the extreme scenario where immediately after production, partons are free streamed until the beginning of the hydrodynamic evolution.
-We note that such an approach has been used in various recent Bayesian analysis, where the most likely range of parameter values for the hydrodynamic scenario is sought Bernhard et al.
-(2019); Everett et al.
-(2021a, b); Parkkila et al.
-(2021); Nijs et al.
-(2021a, b); Liyanage et al.
-(2023).
+We note that such an approach has been used in various recent Bayesian analysis, where the most likely range of parameter values for the hydrodynamic scenario is sought [32, 33, 34, 35, 36, 37, 38].
 We perform our study using well known estimators for elliptic and triangular flows, but we also present estimators that have received little attention for isobars for quantities such as mean transverse momentum and associated cumulants, correlators between elliptic/triangular flow and mean transverse momentum, symmetric cumulant, mixed harmonic correlator.
@@
 The original distribution is spherically symmetric, hence not suitable for studying nuclei which are expected to have a deformed shape, such as ${}^{96}_{40}$r and ${}^{96}_{44}$u .
-The non-spherical shape is then taken into account by introducing an angular dependence in the nuclei radius, such as done in Jia (2022a); Giacalone et al.
-(2021b); Bally et al.
-(2022b).
+The non-spherical shape is then taken into account by introducing an angular dependence in the nuclei radius, such as done in [15, 14, 39].
 Hence, the probability $P(r,\,\theta,\,\varphi)$ of finding a nucleon in the position $\vec{r}=(r,\,\theta,\,\varphi)$ is proportional to the nuclear density profile
@@
 It is expected that variations in these parameters are reflected in final-state observables.
-Here, motivated by the values for these parameters present in KIBÉDI and SPEAR (2002); Pritychenko et al.
-(2016); Zhang and Jia (2022) for ${}^{96}_{44}$u and ${}^{96}_{40}$r , we sample 6 distinct cases, where we progressively tune our parameters from a triaxial ${}^{96}_{44}$u nucleus (case 1) to the ${}^{96}_{40}$r nucleus (case 6).
-The parameters used are shown in Table 1
-{tblr}
-colspec=—Q[c,50pt]—Q[c,50pt]—Q[c,50pt]—Q[c,50pt]—Q[c,50pt]—Q[c,50pt]—, hlines = black, .5pt, vlines = black, .5pt Case Label & $\mathbf{R_{0}}$ (fm) $\mathbf{a}$ (fm) $\mathbf{\beta_{2}}$ $\mathbf{\beta_{3}}$ $\mathbf{\gamma}$
-Case 1 5.09 0.46 0.16 0 $\pi/6$ [Uncaptioned image]
-Case 2 5.09 0.46 0.16 0 0 [Uncaptioned image]
-Case 3 5.09 0.46 0.16 0.20 0 [Uncaptioned image]
-Case 4 5.09 0.46 0.06 0.20 0 [Uncaptioned image]
-Case 5 5.09 0.52 0.06 0.20 0 [Uncaptioned image]
-Case 6 5.02 0.52 0.06 0.20 0 [Uncaptioned image]
+Here, motivated by the values for these parameters present in [40, 41, 18] for ${}^{96}_{44}$u and ${}^{96}_{40}$r , we sample 6 distinct configurations, where we progressively tune our parameters from a triaxial ${}^{96}_{44}$u nucleus (configuration 1) to the ${}^{96}_{40}$r nucleus (configuration 6).
+The parameters used are shown in Table 1.
 Table 1: Nuclear structure parameters employed for generating initial conditions of isobar collisions.
+Configuration Label $\mathbf{R_{0}}$ (fm) $\mathbf{a}$ (fm) $\mathbf{\hskip 5.69046pt\beta_{2}\hskip 5.69046pt}$ $\mathbf{\hskip 5.69046pt\beta_{3}\hskip 5.69046pt}$ $\mathbf{\gamma}$
+Configuration 1 5.09 0.46 0.16 0 $\pi/6$
+Configuration 2 5.09 0.46 0.16 0 0
+Configuration 3 5.09 0.46 0.16 0.20 0
+Configuration 4 5.09 0.46 0.06 0.20 0
+Configuration 5 5.09 0.52 0.06 0.20 0
+Configuration 6 5.02 0.52 0.06 0.20 0
 In Table 1, when going from one line to the next, only one parameter is changed.
-Cases 1 and 2 are similar except that the former is a triaxial ellipsoid and the later is a spheroid.
-An octupole deformation (pear shape) has been introduced when going from case 2 to 3.
+Configurations 1 and 2 are similar except that the former is a triaxial ellipsoid and the later is a spheroid.
+An octupole deformation (pear shape) has been introduced when going from configuration 2 to 3.
 When going from case 3 to 4, the quadrupole deformation (ellipsoid shape) is reduced.
-Cases 4 and 5 are similar except that the later has a larger diffuseness.
-Lastly, the nuclear radius is reduced when going from case 5 to 6.
-These systematic modifications in the nuclear structure parameters allow to investigate eventual changes in the results when gradually moving from what is typically considered a Ruthenium nucleus (case 2) to what typically is used as a Zirconium nucleus (case 6).
+Configurations 4 and 5 are similar except that the later has a larger diffuseness.
+Lastly, the nuclear radius is reduced when going from configuration 5 to 6.
+These systematic modifications in the nuclear structure parameters allow to investigate eventual changes in the results when gradually moving from what is typically considered a Ruthenium nucleus (configuration 2) to what typically is used as a Zirconium nucleus (configuration 6).
+The effect of these different parameters in the shape of the nucleus can be seen in Fig.
+1.
 Effects related to “neutron-skin” have not been included in our simulations, therefore, protons and neutrons follow the same nuclear distribution.
-The reader may refer to Xu et al.
-(2021b) for a study on the neutron-skin effects.
-Due to its flexibility and the ability to accept pre-generated (possibly deformed) nuclear configurations for projectile and target, the TRENTo model Moreland et al.
-(2015) has been used to generate initial conditions for our simulations.
-Two nuclear configurations (1) are randomly selected from a set of 50.000 pre-generated options and subsequently randomized in orientation.
+The reader may refer to [42] for a study on the neutron-skin effects.
+Figure 1: Shapes of different configurations in the nuclei, using the parameters of Table 1 in Eq. (1).
+Due to its flexibility and the ability to accept pre-generated (possibly deformed) nuclear configurations for projectile and target, the TRENTo model [43] has been used to generate initial conditions for our simulations.
+Two nuclear configurations (1) are randomly selected from a set of $50000$ pre-generated options and subsequently randomized in orientation.
 The Monte-Carlo Glauber model, where a value for the nucleon-nucleon cross-section $\sigma_{\rm NN}$ is needed, is then applied to identify which nucleons will participate in the collision, and to each participating nucleon, a density function $\rho_{i}(x,y,z)$ is assigned, namely a Gaussian function with standard deviation given by the nucleon width $w$.
@@
 These thicknesses are then used to calculate the energy density $\epsilon(x,y,\eta_{s}=0)$ as a reduced thickness between the two participant thicknesses given by
-$\lim_{\tau\to 0^{+}}\tau\epsilon(x,y,\eta_{s}=0)=N\left[\frac{T_{A}(x,y)+T_{B}(x,y)}{2}\right]^{1/p}\,,$ (2)
+$\bar{\epsilon}(x,y)\equiv\lim_{\tau\to 0^{+}}\tau\epsilon(x,y,\eta_{s}=0)=N\left[\frac{T_{A}^{p}(x,y)+T_{B}^{p}(x,y)}{2}\right]^{1/p}\,,$ (2)
 where $N$ is a normalization constant and $p$ is called the reduced thickness parameter.
@@
 However, in order to actually produce initial conditions with TRENTo , one needs to specify a value for its physical parameters ($N$, $w$, $\sigma_{\rm NN}$, $k$, $p$).
-In our simulations, these parameters were chosen to be the same as in Everett et al.
-(2021b) (because their analysis was done for both RHIC and LHC energies) and are outlined in Table 2 for completeness.
+In our simulations, these parameters were chosen to be the same as in [34] (because their analysis was done for both RHIC and LHC energies) and are outlined in Table 2 for completeness.
+Table 2: TRENTo parameters used to generate initial conditions[34]
 Parameter Symbol Value
@@
 Reduced thickness parameter $p$ 0.063
-Table 2: TRENTo parameters used to generate initial conditionsEverett et al. (2021b).
-After the collision takes place, one needs to specify how long the produced system will stay in the prethermalization phase.
+After the collision takes place, one needs to specify how long the produced system will stay in the pre-thermalization stage.
 Given the model chosen in this work, such characteristic is controlled by the free-streaming time, $\tau_{FS}$.
-It is worth noting that at the moment all assumptions related to the modeling of pre-equilibrium phase are ad-hoc and such a quantity is currently chosen to have either a constant value or is allowed to carry a dependence with the collision centrality.
-Regardless of the case, the value of $\tau_{FS}$ or its behavior with the collision centrality started to be determined (together with many other parameters needed in a complete hybrid hydrodynamic simulation) via Bayesian inference techniques Moreland et al.
-(2020); Bernhard et al.
-(2019); Everett et al.
-(2021b, a); Nijs et al.
-(2021a); Parkkila et al.
-(2021); Nijs and van der Schee (2022); Parkkila et al.
-(2022).
-This does not mean, however, that the main characteristics of this phase are well understood.
-In fact, it depends on the many particularities of each statistical analysis as different collaborations have obtained values for $\tau_{FS}$ that vary more than 50%: for instance, Parkkila et al.
-(2021) extracted $\tau_{FS}=0.901$ fm/c by considering Pb+Pb collisions at 5.02 TeV; on the other hand, Parkkila et al.
-(2022) reported $\tau_{FS}=0.71$ fm/c improving on the previous study by considering a simultaneous analysis of Pb+Pb data at 2.76 TeV and 5.02 TeV.
-When allowing the free-streaming to change with centrality as in Everett et al.
-(2021b), with parameters from the Grad viscous correction model, for Au+Au collisions at 200 GeV, $1.36\lesssim\tau_{FS}\,({\rm fm/c})\lesssim 1.51$, a variation of almost 50% when compared to the improved estimate from Parkkila et al.
-(2022).
-Lastly, a simultaneous analysis of p+Pb and Pb+Pb data at 5.02 TeV presented in Moreland et al.
-(2020) yielded $\tau_{FS}=0.37$ fm/c.
-While a smaller free-streaming time is expected for p+A collisions due to the system size after the collision takes place, this estimate (which is also applied to Pb+Pb collisions if one makes use of the model outlined in Moreland et al.
-(2020)) is $\sim 0.59\%$ smaller than the value quoted in Parkkila et al.
-(2021) which only considered Pb+Pb data at the same collision energy.
-Given the current lack of a detailed understanding of the prethermalization phase and the large uncertainty in the extracted values for $\tau_{FS}$, in the present study we consider three different choices for the free-streaming time: (i) $\tau_{FS}=0.4$ fm/c, (ii) $\tau_{FS}=1.0$ fm/c and (iii) a centrality-dependent free-streaming time as implemented in Everett et al.
-(2021b): $\tau_{FS}=\tau_{R}(\{\epsilon\}/\{\epsilon_{R}\})^{\alpha}$, where $\{\epsilon\}$ denotes a (particular) average initial energy density in the transverse plane for a given collision, $\{\epsilon_{R}\}=4.0$ GeV/fm2, $\tau_{R}=1.46$ fm/c and $\alpha=0.031$ are constants that were adjusted by comparison with experimental data.
-These choices cover the range of values determined by recent Bayesian analysis and allow for a systematic investigation of eventual impacts that the prethermalization phase may have on quantities related to the system shape.
+It is worth noting that at the moment all assumptions related to the modeling of pre-equilibrium stage are ad-hoc and such a quantity is currently chosen to have either a constant value or is allowed to carry a dependence with the collision centrality.
+Regardless of the case, the value of $\tau_{FS}$ or its behavior with the collision centrality started to be determined (together with many other parameters needed in a complete hybrid hydrodynamic simulation) via Bayesian inference techniques [44, 32, 34, 33, 36, 35, 45, 46].
+This does not mean, however, that the main characteristics of this stage are well understood.
+In fact, it depends on the many particularities of each statistical analysis as different collaborations have obtained values for $\tau_{FS}$ that vary more than 50%: for instance, [35] extracted $\tau_{FS}=0.901$ fm/c by considering Pb+Pb collisions at 5.02 TeV; on the other hand, [46] reported $\tau_{FS}=0.71$ fm/c improving on the previous study by considering a simultaneous analysis of Pb+Pb data at 2.76 TeV and 5.02 TeV.
+When allowing the free-streaming to change with centrality as in [34], with parameters from the Grad viscous correction model, for Au+Au collisions at 200 GeV, $1.36\lesssim\tau_{FS}\,({\rm fm/c})\lesssim 1.51$, a variation of almost 50% when compared to the improved estimate from [46].
+Lastly, a simultaneous analysis of p+Pb and Pb+Pb data at 5.02 TeV presented in [44] yielded $\tau_{FS}=0.37$ fm/c.
+While a smaller free-streaming time is expected for p+A collisions due to the system size after the collision takes place, this estimate (which is also applied to Pb+Pb collisions if one makes use of the model outlined in [44]) is $\approx 0.59\%$ smaller than the value quoted in [35] which only considered Pb+Pb data at the same collision energy.
+Given the current lack of a detailed understanding of the pre-thermalization stage and the large uncertainty in the extracted values for $\tau_{FS}$, in the present study we consider three different choices for the free-streaming time: (i) $\tau_{FS}=0.4$ fm/c, (ii) $\tau_{FS}=1.0$ fm/c and (iii) a centrality-dependent free-streaming time as implemented in [34]: $\tau_{FS}=\tau_{R}(\{\bar{\epsilon}\}/\bar{\epsilon}_{R})^{\alpha}$, where $\{{\bar{\epsilon}}\}$ denotes a (particular) average initial energy density in the transverse plane for a given collision, $\bar{\epsilon}_{R}=4.0$ GeV/fm2, $\tau_{R}=1.46$ fm/c and $\alpha=0.031$ are constants that were adjusted by comparison with experimental data.
+These choices cover the range of values determined by recent Bayesian analysis and allow for a systematic investigation of eventual impacts that the pre-thermalization stage may have on quantities related to the system shape.
 In Fig.
-1, the free-streaming time for the centrality-dependent case is shown as function of $\{\epsilon\}$ for Ru (very similar results hold for Zr).
+2, the free-streaming time for the centrality-dependent case is shown as function of $\{{\bar{\epsilon}}\}$ for Ru (very similar results hold for Zr).
 Centrality windows were estimated from ordering the total energy $E$ (integral on the transverse plane of the energy density in eq.
-(2) of 1 million TRENTo initial conditions in descending order for each colliding system.
+(2), $E=N\int d^{2}x_{\perp}\left[\frac{T_{A}^{p}(x,y)+T_{B}^{p}(x,y)}{2}\right]^{1/p}$) of 1 million TRENTo initial conditions in descending order for each colliding system.
 One sees that the free-streaming time varies between $\approx$ 1.35 and $\approx$ 1.49 fm in such a case.
-Figure 1: Free-streaming time as function of $\{\epsilon\}$ in a mid-rapidity slice with $\tau_{R}=1.46$ fm/c, $\{\epsilon_{R}\}=4$ GeV/fm2 and $\alpha=0.031$.
-Centralities windows are also shown (see text).
+In Fig.
+2, $\tau_{FS}$, the time when free-streaming stops and hydrodynamics is switched on, increases for larger energy densities.
+Though this is discussed and some arguments for this put forward in [34], one might expect the opposite: for larger densities, more collisions happen and hydrodynamics starts earlier.
+Therefore we have also included a curve that corresponds to a negative value for $\alpha$.
+For central collisions, which are usually considered more interesting for isobar collisions, both positive and negative $\alpha$ lead to similar fairly constant values.
+However we will discuss both cases when necessary below.
+Figure 2: Free-streaming time as function of $\{{\bar{\epsilon}}\}$ in a mid-rapidity slice with $\tau_{R}=1.46$ fm/c, $\bar{\epsilon}_{R}=4$ GeV/fm2 and $\alpha=0.031$ [34].
+Centrality windows are also shown (see text).
 Very similar results are obtained for ${{}_{40}^{96}Zr}+{{}_{40}^{96}Zr}$.
+The $\alpha=-0.031$ case is shown by the red dot-dashed line.
+Both positive and negative $\alpha$’s lead to comparable values of $\tau_{FS}$ for central collisions.
 III Interplay between free-streaming and nuclear geometry
@@
 $\varepsilon_{n}\{4\}$ =$4]{2\langle\varepsilon_{n}^{2}\rangle^{2}-\langle\varepsilon_{n}^{4}\rangle}\,.$ (5)
+Eccentricities have long been expected to be good estimators for anisotropic flow coefficients [47, 24, 48].
+It has been argued that when there is a free-streaming stage and a Laudau matching is done to switch to a hydrodynamic description, an initial momentum anisotropy is created and could also play a role [49, 50].
+More recent papers indicate that for nuclear collisions (and some pre-equilibrium stage), initial spatial eccentricity is a good predictor for elliptic flow (see e.g.
+[51, 52]).
+Therefore in this paper, we concentrate on spatial eccentricities.
 Fig.
-2 shows the same event at the initial time and at the end of the free-streaming.
+3 shows the same event at the initial time and at the end of the free-streaming (computed for configuration (iii)).
 One sees that the energy density distribution becomes more rounded, with a larger average radius $R\equiv\sqrt{\{r^{2}\}}$ as time flows.
@@
 This is shown more precisely in Fig.
-3.
+4.
 In this figure, the ratio between a given quantity computed after free-streaming and for the initial condition is used.
-It is a more effective mean of expressing the influence of the prethermalization phase.
+It is a more effective mean of expressing the influence of the pre-thermalization stage.
 The following concise notation is used
 ${\rm ratio}[A]=\frac{\langle A\rangle_{FS}}{\langle A\rangle_{IC}}.$
-Here and below, to determine centrality bin edges, all generated events are used, regardless of the (geometrical) case they belong to in Table 2.
-Such a procedure is different from the actual experiment, where centrality classification is based on the number of charged tracks together with information from Monte Carlo Glauber simulations for each collision system independently Abdallah et al.
-(2022).
-Results are in general shown in bins of width 1% in the 0-5% centrality window and then in bins of 5%; finer bins are used to $\rho_{2}$ and $\rho_{3}$, see below.
-When statistics demands, the 5 most central bins are merged in a single one of 0-5%.
+Here and below, to determine centrality bin edges, all generated events are used, regardless of the (geometrical) configuration they belong to in Table 2.
+Such a procedure is different from the actual experiment, where centrality classification is based on the number of charged tracks together with information from Monte Carlo Glauber simulations for each collision system independently [2].
+Results are in general shown in bins of width 1% in the 0–5% centrality window and then in bins of 5%; finer bins are used to $\rho_{2}$ and $\rho_{3}$, see below.
+When statistics demands, the 5 most central bins are merged in a single one of 0–5%.
 Centrality bin edges can be read from Table 3
+Table 3: Definition of the centrality bins as a function of E.
 Centrality Bin Energy Range (GeV) Centrality Bin Energy Range (GeV)
-0-1% 700.00 - 466.10 10-15% 275.69 - 225.49
-1-2% 466.10 - 444.98 15-20% 225.49 - 183.31
-2-3% 444.98 - 426.64 20-25% 183.31 - 147.75
-3-4% 426.64 - 409.73 25-30% 147.75 - 117.71
-4-5% 409.73 - 336.08 30-35% 117.71 - 92.53
-5-10% 336.08 - 275.69 35-40% 92.53 - 71.62
-Table 3: Definition of the centrality bins in function of E.
+0 – 1% 700.00 – 466.10 10-15% 275.69 – 225.49
+1 – 2% 466.10 – 444.98 15-20% 225.49 – 183.31
+2 – 3% 444.98 – 426.64 20-25% 183.31 – 147.75
+3 – 4% 426.64 – 409.73 25-30% 147.75 – 117.71
+4 – 5% 409.73 – 336.08 30-35% 117.71 – 92.53
+5 – 10% 336.08 – 275.69 35-40% 92.53 – 71.62
 In Fig.
-3, curves end up grouped by how long the produced system free-streamed: for eccentricities, one group of curves corresponds to $\tau_{FS}=0.4$ fm/c, the middle one to $\tau_{FS}=1.0$ fm/c and the last one to the centrality-dependent $\tau_{FS}$, estimated to be approximately between 1.35 fm/c and 1.49 fm/c from Fig.
-1.
-Free-streaming effects are increasingly stronger (meaning, larger deviations from 1) going from $R$, to $\epsilon_{2}$, then $\epsilon_{3}$ and finally $\epsilon_{4}$.
-Ratios for the 6 different geometrical configurations exhibit extremely similar shapes which mean they are affected fairly similarly by the free-streaming modeling of the prethermalization phase.
-A conformal expansion during the free-streaming is assumed.
-Since this model introduces no interactions as time goes by, in the case of a free-streaming velocity smaller than $c$ the system would have a slower expansion, reducing the effects of the prethermalization phase.
-An interacting model will introduce correlations on the system that will further slow down the eccentricity reductions.
+4, curves end up grouped by how long the produced system free-streamed: for eccentricities, one group of curves corresponds to $\tau_{FS}=0.4$ fm/c, the middle one to $\tau_{FS}=1.0$ fm/c and the last one to the centrality-dependent $\tau_{FS}$, estimated to be approximately between 1.35 fm/c and 1.49 fm/c from Fig.
+2.
+Free-streaming effects are increasingly stronger (meaning, larger deviations from 1) going from $R$, to $\varepsilon_{2}$, then $\varepsilon_{3}$ and finally $\varepsilon_{4}$.
+Ratios for the 6 different geometrical configurations exhibit extremely similar shapes (which mean they are affected fairly similarly by the free-streaming modeling of the pre-thermalization stage) so results are only shown for configurations 2 and 6.
+In the approach we use, the free-streaming velocity is equal to $c$.
+In case it is smaller, the system would have a slower expansion, reducing the effects of the pre-thermalization stage.
+An interacting model will introduce correlations on the system that will also attenuate the effects of the pre-thermalization stage.
 In this way, the magnitude of the effects shown here may be understood as an upper bound.
-For the subsequent hydrodynamic phase, reducing the free-streaming velocity is important for mitigating the artificial increase of the bulk pressure at the start of the expansion da Silva et al.
-(2023).
-Figure 2: The mid-rapidity energy density slice of a ${}^{96}_{40}$r +${}^{96}_{40}$r event in the 20-30% centrality window is shown (left) initially and (right) after free-streaming.
+For the subsequent hydrodynamic phase, reducing the free-streaming velocity is important for mitigating the artificial increase of the bulk pressure at the start of the expansion [53].
+Figure 3: The mid-rapidity energy density slice of a ${}^{96}_{40}$r +${}^{96}_{40}$r event in the 20–30% centrality window is shown (left) initially and (right) after free-streaming.
 We now concentrate on effects of varying the geometry, for a fixed free-streaming time.
@@
 Fig.
-3) and ratios for a given configuration 1-5 with respect to 6 are essentially equal to the ones obtained at initial condition level, so not shown.
-Unless otherwise stated, we ran 10 million events minimum-bias to obtain the results shown.
-Figure 3: Root mean square radius $R$ and eccentricities $\epsilon_{n}$ at the end of free-streaming phase divided by their initial values.
+4) and ratios for a given configuration 1–5 with respect to 6 are essentially equal to the ones obtained at initial condition level, so not shown.
+Unless otherwise stated, we ran 10 million events minimum-bias for each nuclear configuration present in Table 1 to obtain the results shown.
+Figure 4: Root mean square radius $R$ and eccentricities $\varepsilon_{n}$ at the end of free-streaming stage divided by their initial values.
 Line styles correspond to different free-streaming times.
 One million minimum-bias events were used to generate this figure.
-Figure 4: Left panels: Ratios between ellipticity computed for cases 1 through 5 and case 6, as function of total initial energy and with different free-streaming times.
+Results for all configurations lie close so only those for configurations 2 and 6 are shown.
+Figure 5: Left panels: Ratios between ellipticity computed for configurations 1 through 5 and configuration 6, as function of total initial energy and with different free-streaming times.
 From top to bottom, the calculation is made for mean ellipticity, second and fourth order cumulants.
@@
 Configuration 3 shows a higher ellipticity than configuration 2 even though both configurations only differ by the inclusion of an octupole deformation $\beta_{3}$, indicating the non-trivial interplay between $\beta_{2}$ and $\beta_{3}$ discussed e.g.
-in Jia (2022a).
+in [15].
 The decrease in ellipticity from configuration 3 to 4 is expected due to the decrease of the quadrupole deformation.
@@
 The differences in geometry are more apparent for the most central collisions.
-For the second order cumulant ratios: $\epsilon_{2}\{2\}_{n}/\epsilon_{2}\{2\}_{6}=\langle\epsilon_{2}\rangle_{n}/\langle\epsilon_{2}\rangle_{6}\times\sqrt{1+(\sigma^{2}_{n}/\langle\epsilon_{2}\rangle^{2}_{n})}^{2}\\ /\sqrt{1+(\sigma^{2}_{6}/\langle\epsilon_{2}\rangle^{2}_{6})^{2}}$.
-We have checked that the quantities $\sigma^{2}_{n}/\langle\epsilon_{2}\rangle^{2}_{n}$ are nearly equal for n=1-6, so it explains the similarity between the upper and lower panels.
-The ratios of $\epsilon_{2}\{4\}$ are also shown, even though the error bars are large, because it has been argued that the comparison of $v_{2}\{2\}$ and $v_{2}\{4\}$ or $\epsilon_{2}\{2\}$ and $\epsilon_{2}\{4\}$ for isobars might yield information on nuclear structure Jia et al.
-(2023).
+The ratio $\langle\varepsilon_{2}\rangle_{i}/\langle\varepsilon_{2}\rangle_{6}$ is very close to the ratio $\varepsilon_{2}\{2\}_{i}/\varepsilon_{2}\{2\}_{6}$ so only the latter is shown.
+This is because $\varepsilon_{2}\{2\}_{i}/\varepsilon_{2}\{2\}_{6}=\langle\varepsilon_{2}\rangle_{i}/\langle\varepsilon_{2}\rangle_{6}\times\sqrt{1+(\sigma^{2}_{i}/\langle\varepsilon_{2}\rangle^{2}_{i})}^{2}/\sqrt{1+(\sigma^{2}_{6}/\langle\varepsilon_{2}\rangle^{2}_{6})^{2}}$ and we have checked that the quantities $\sigma^{2}_{n}/\langle\varepsilon_{2}\rangle^{2}_{n}$ are nearly equal for $i=1-6$.
+The ratios of $\varepsilon_{2}\{4\}$ are also shown, even though the error bars are large, because it has been argued that the comparison of $v_{2}\{2\}$ and $v_{2}\{4\}$ or $\varepsilon_{2}\{2\}$ and $\varepsilon_{2}\{4\}$ for isobars might yield information on nuclear structure [54].
 The inset illustrates the effect of changing the diffuseness $a$.
@@
 Other details such as the nuclear radius, triaxiality, and quadrupole deformation do not affect this quantity.
-From the right hand side, it can be seen that free-streaming has larger effects for triangularity than for ellipticity for cases 1 and 2 (divided by that of case 6): there is a larger splitting for the two different free-streaming times, almost 1% for most centralities, which is comparable with the error bars in experimental triangular flow ratios.
-For cases 3-5, the effect is a little less pronounced.
-In any case, the modeling and duration of the prethermalization phase may play a role in a very precise determination of the octopole deformation.
+From the right hand side, it can be seen that free-streaming has larger effects for triangularity than for ellipticity for configurations 1 and 2 (divided by that of configuration 6): there is a larger splitting for the two different free-streaming times, almost 1% for most centralities, which is comparable with the error bars in experimental triangular flow ratios.
+For configurations 3–5, the effect is a little less pronounced.
+In any case, the modeling and duration of the pre-thermalization stage may play a role in a very precise determination of the octopole deformation.
 Figure 6: Ratios of triangularities, similar to Fig. 5.
 Figure 7 shows a higher-order eccentricity, $\varepsilon_{4}$, as function of the total initial energy $E$ for different free-streaming times.
-When going from case 4 to case 5, for central collisions, the dependence on the diffuseness parameter is 3% for central collisions and larger for non central ones.
+When going from configuration 4 to configuration 5, for central collisions, the dependence on the diffuseness parameter is 3% for central collisions and larger for non central ones.
 In line with results obtained for triangularity, a progressively larger splitting between the results with different free-streaming times is observed.
@@
 Figure 8: Ratios of root mean square of radius, in a similar fashion as done in Fig. 5 for ellipticity.
-Given that a good mapping is expected between not just $v_{2}$, $v_{3}$ and $\epsilon_{2}$, $\epsilon_{3}$, but also between the event plane angles and initial state symmetry plane angles $\Phi_{2}$, $\Phi_{3}$, it can be used to study the normalized symmetric cumulant NSC(2,3) Gardim et al.
-(2017),
+Given that a good mapping is expected between not just $v_{2}$, $v_{3}$ and $\varepsilon_{2}$, $\varepsilon_{3}$, but also between the event plane angles and initial state symmetry plane angles $\Phi_{2}$, $\Phi_{3}$, it can be used to study the normalized symmetric cumulant $\varepsilon NSC(2,3)$ [55],
 $\varepsilon NSC(2,3)=\frac{\langle\varepsilon_{2}^{2}\varepsilon_{3}^{2}\rangle-\langle\varepsilon_{2}^{2}\rangle\langle\varepsilon_{3}^{2}\rangle}{\langle\varepsilon_{2}^{2}\rangle\langle\varepsilon_{3}^{2}\rangle}\,$ (6)
 as well as the two-plane correlator
-$\langle\cos 6(\Phi_{2}-\Phi_{3})\rangle=\frac{\langle\epsilon_{2}^{3}\epsilon_{3}^{2}\cos 6(\Phi_{2}-\Phi_{3})\rangle}{\sqrt{\langle\epsilon_{2}^{6}\rangle\langle\epsilon_{3}^{4}\rangle}}\,.$ (7)
-Figure 8: Normalized symmetric cumulant $\epsilon NSC(2,3)$ computed from initial conditions for the 6 geometrical cases.
+$\langle\cos 6(\Phi_{2}-\Phi_{3})\rangle=\frac{\langle\varepsilon_{2}^{3}\varepsilon_{3}^{2}\cos 6(\Phi_{2}-\Phi_{3})\rangle}{\sqrt{\langle\varepsilon_{2}^{6}\rangle\langle\varepsilon_{3}^{4}\rangle}}\,.$ (7)
+Figure 9: Ratio of normalized symmetric cumulant $\varepsilon NSC(2,3)$ computed for the geometrical configurations $i=1-5$ compared to $i=6$.
 In Fig.
-(8), for the normalized symmetric cumulant computed directly from the initial conditions, cases 1, 2, 3 are similar and 4 and 5 are fairly equal.
+9, for the normalized symmetric cumulant computed directly from the initial conditions, configurations 1, 2, 3 are similar and 4 and 5 are fairly equal and are shown separately for clarity.
 Free streaming seems to have a strong effect, and might change once or twice the sign of the ratio, however error bars are large, so more statistics would be necessary to confirm this.
+There are oscillations for $E<200$ GeV so results are shown only for larger values.
 The changes of sign can be traced back to slightly different free-streaming time dependencies for each nuclear geometry considered: as the system evolves in time, the individual quantities $\varepsilon NSC(2,3)$ appearing in this ratio becomes less negative, with each of them crossing zero at different energy values.
-Case 6 only shows a single change of sign (positive to negative) that always happens for values of energy lower than the ones shown in our plots, meanwhile, cases 2 and 3 for instance, present two clear changes of sign (positive to negative and back to positive), with the second change happening in the most central bin considered.
+Configuration 6 only shows a single change of sign (positive to negative) that always happens for values of energy lower than the ones shown in our plots, meanwhile, configurations 2 and 3 for instance, present two clear changes of sign (positive to negative and back to positive), with the second change happening in the most central bin considered.
+At this point, it is interesting to discuss what would happen if $\alpha$ were negative in the definition of $\tau_{FS}$ in case (iii).
+In the leftmost bin shown in the figure, according to fig.
+2, the values of $\tau_{FS}$ are fairly equal so the ratio of $\varepsilon NSC(2,3)$ is the same for $\alpha$ positive or negative.
+In the rightmost bin, again from fig.
+2, $\tau_{FS}$ would decrease from $\approx 1.48$ fm/c to $\approx 1.44$ fm/c.
+Therefore the value of the ratio of $\varepsilon NSC(2,3)$ would be between the value for variable $\tau_{FS}$ and that for $\tau_{FS}=1$ fm/c, but nearest to the former.
+So the sign changes (top panel) would not be much affected.
 The ratio of two-plane correlators, fluctuates around zero and suffer from lack of statistics so is not shown.
 Another interesting angle correlator for which data have been obtained is $ac2(3)=\langle v_{2}^{2}v_{4}\cos 4(\Psi_{2}-\Psi_{4})\rangle$.
-Since it cannot be predicted for all centralities with just $\varepsilon_{4}$ Gardim et al.
-(2012b, 2015b), we leave it to a future study from hydrodynamics.
-We note however that it has a good potential to discriminate geometry in central collisions Zhao et al.
-(2023b).
+Since it cannot be predicted for all centralities with just $\varepsilon_{4}$ [47, 24], we leave it to a future study from hydrodynamics.
+We note however that it has a good potential to discriminate geometry in central collisions [56].
 The observables presented so far, except perhaps $\varepsilon NSC(2,3)$, exhibit minimal sensitivity to the effects of free-streaming and triaxiality.
-One observable known to be sensitive to triaxiality is the Pearson correlator between event anisotropic flow and event mean transverse momentum $[p_{T}]$ in a given centrality bin Bally et al.
-(2022a); Aad et al.
-(2023).
-This correlator has been identified as a strong candidate for assessing the effects on triaxility in isobars Jia (2022b).
+One observable known to be sensitive to triaxiality is the Pearson correlator between event anisotropic flow and event mean transverse momentum $[p_{T}]$ in a given centrality bin [21, 8].
+This correlator has been identified as a strong candidate for assessing the effects on triaxility in isobars [16].
 Previous studies were limited to $n=2$ and utilized estimators that lack proper correlation with mean transverse momentum: Ref.
-Giacalone et al.
-(2021c) showed that a good proxy to estimate this type of correlator is to replace the mean momentum by the ratio between total energy and total entropy in a very narrow centrality bin and both $v_{2}\propto\varepsilon_{2}$ and $v_{3}\propto\varepsilon_{3}$ can be considered.
+[52] showed that a good proxy to estimate this type of correlator is to replace the mean momentum by the ratio between total energy and total entropy in a very narrow centrality bin and both $v_{2}\propto\varepsilon_{2}$ and $v_{3}\propto\varepsilon_{3}$ can be considered.
 Therefore we calculate the following Pearson correlator:
@@
 where the averages are taken in centrality bins of 0.25% centrality and $\sigma_{\varepsilon_{n}^{2}}$ and $\sigma_{E/S}$ are the standard deviation of $\varepsilon_{n}$ and $E/S$ in those centrality bins, respectively.
+$S$ is computed as $S=\int d^{2}x_{\perp}s\left(N\left[\frac{T_{A}(x,y)^{p}+T_{B}^{p}(x,y)}{2}\right]^{1/p}\right)$ where the entropy density $s$ is obtained from the equation of state.
 […]
 Our results confirm that the effect the triaxiality parameter $\gamma$ is particularly clear when comparing configurations 1 and 2 for $\rho_{2}$ in the top panel of Fig.
-9.
-We also report that the splitting between the results of these two cases increases with the free-streaming time, starting from $\sim 0.3$ at the initial condition level and reaching $\sim 0.5$ when considering a centrality dependent free-streaming duration.
-Comparing cases 2 and 3, shows that the value of $\beta_{3}$ has a strong effect, lowering the correlation in the $\sim 5\%-13\%$ centrality range.
-Moreover, one verifies that the free-streaming time plays an interesting role, inducing a single or a double sign change depending on how long the system stays in the prethermalization phase.
+10.
+We also report that the splitting between the results of these two configurations increases with the free-streaming time, starting from $\approx 0.3$ at the initial condition level and reaching $\approx 0.5$ when considering a centrality dependent free-streaming duration.
+Comparing configurations 2 and 3, shows that the value of $\beta_{3}$ has a strong effect, lowering the correlation in the $\approx 5\%-13\%$ centrality range.
+Moreover, one verifies that the free-streaming time plays an interesting role, inducing a single or a double sign change depending on how long the system stays in the pre-thermalization stage.
 While we still lack final experimental results for this correlator, such a signature may eventually help to constrain the duration of the pre-hydrodynamic phase in isobar collisions.
 Next, reducing the quadrupole deformation while keeping the octupole deformation fixed shifts the correlation to values around unity with a much reduced free-streaming time dependence.
-For the case 5, the result is brought even closer to unity.
+For the configuration 5, the result is brought even closer to unity.
 Using these results together with the ones presented in Fig.
-4, 5 indicates that it might be possible to constrain $\beta_{2}$, $\beta_{3}$ and $\gamma$ simultaneously from observations of anisotropic flow and the correlation of elliptical flow with mean transverse momentum.
-We have checked that using different estimators for $[p_{T}]$, the effect of the free-streaming time is still substantial.
-Moving to $\rho_{3}$, and assuming that the estimator (8) is still a good proxy for the final state observable (as is the case for high-energy Pb+Pb collisions Giacalone et al.
-(2021c)), one sees that while the dependence on the triaxiality parameter becomes weaker, the effect of free-streaming time is still visible for cases 1-3.
-In addition, adding an octupole deformation on top of a quadrupole deformation (case 3) shifts the correlation to even larger values in central collisions, providing a clear separation with cases 1 and 2 where only the quadrupole deformation is present.
-Cases 4, 5 and 6 all have similar $\rho_{3}$, indicating that this observable is independent of diffuseness and distribution radius in the centrality range considered.
+5, 6 indicates that it might be possible to constrain $\beta_{2}$, $\beta_{3}$ and $\gamma$ simultaneously from observations of anisotropic flow and the correlation of elliptical flow with mean transverse momentum.
+We have checked that using different estimators for $[p_{T}]$ (taken from  [52]), the effect of the free-streaming time is still substantial.
+Moving to $\rho_{3}$, and assuming that the estimator (8) is still a good proxy for the final state observable (as is the case for high-energy Pb+Pb collisions [52]), one sees that while the dependence on the triaxiality parameter becomes weaker, the effect of free-streaming time is still visible for configurations 1–3.
+In addition, adding an octupole deformation on top of a quadrupole deformation (configuration 3) shifts the correlation to even larger values in central collisions, providing a clear separation with configurations 1 and 2 where only the quadrupole deformation is present.
+Configurations 4, 5 and 6 all have similar $\rho_{3}$, indicating that this observable is independent of diffuseness and distribution radius in the centrality range considered.
 Figure 10: Top panel: Pearson correlation between eccentricity squared and the ratio of total energy by total entropy as function of total initial energy for initial conditions.
 Lower panel: The same, but replacing ellipticity by triangularity in the $\rho_{n}$ correlator.
-Figure 10: Expected value for the distribution of E/S (in a centrality bin) for case n divided by that of case 6 as function of centrality.
+Figure 11: Expected value for the distribution of $E/S$ (in a centrality bin) for configuration $i$ divided by that of configuration 6 as function of centrality.
+The $\alpha<0$ case can be studied in a similar way as done for Fig. 9, one sees that it would not alter the general behavior of $\rho_{2}$ and $\rho_{3}$.
 Another possibility to determine the geometry is the study of the distribution of event mean transverse momentum in a given centrality bin, $[p_{T}]$.
@@
 It is expected that a good estimator for $\langle\delta[p_{T}]^{p}\rangle/\langle[p_{T}]\rangle^{p}$ would be $\langle(\delta(E/S))^{p}\rangle/\langle E/S\rangle^{p}$ where $\delta O\equiv O-\langle O\rangle$ (as far as we know this has not been tested throughout).
-The ratio $\langle E/S\rangle_{n}/\langle E/S\rangle_{6}$ as function of centrality is shown in Fig.
-10.
-It is computed directly from the initial conditions (as E/S is conserved) and exhibit a clear ordering: cases n=1-2 have the highest values for the ratio, above 1, and do not depend on the triaxiality.
-Case 3 is also above 1 but introducing $\beta_{3}$ has decreased the ratio.
-For case 4, decreasing $\beta_{2}$ increases the ratio.
-For case 5, increasing $a$ decreases the ratio.
+The ratio $\langle E/S\rangle_{i}/\langle E/S\rangle_{6}$ as function of centrality is shown in Fig.
+11.
+It is computed directly from the initial conditions (as E/S is conserved) and exhibit a clear ordering: configurations $i=1-2$ have the highest values for the ratio, above 1, and do not depend on the triaxiality.
+Configuration 3 is also above 1 but introducing $\beta_{3}$ has decreased the ratio.
+For configuration 4, decreasing $\beta_{2}$ increases the ratio.
+For configuration 5, increasing $a$ decreases the ratio.
 However, all these effects are quite small.
 Fig.
-11 is more interesting: the ratios $\left(\langle(\delta(E/S))^{2}\rangle/\langle E/S\rangle^{2}\right)_{n}/\left(\langle(\delta(E/S))^{2}\rangle/\langle E/S\rangle^{2}\right)_{6}$ are shown and larger effects are observed.
-Cases n=1-2 have the most extreme values for the ratio, below 0.98 and above 1.01, and do not depend on the triaxiality.
-Case 3 also reaches above 1.01 but introducing $\beta_{3}$ has pushed the ratio towards 1 in general (by about 2%).
-For case 4, decreasing $\beta_{2}$ decreases the ratio.
-For case 5, increasing $a$ pushes the ratio very near 1 except for very central collisions.
+12 is more interesting: the ratios $\left[\langle(\delta(E/S))^{2}\rangle/\langle E/S\rangle^{2}\right]_{i}/\left[\langle(\delta(E/S))^{2}\rangle/\langle E/S\rangle^{2}\right]_{6}$ are shown and larger effects are observed.
+Configurations $i=1-2$ have the most extreme values for the ratio, below 0.98 and above 1.01, and do not depend on the triaxiality.
+Configuration 3 also reaches above 1.01 but introducing $\beta_{3}$ has pushed the ratio towards 1 in general (by about 2%).
+For configuration 4, decreasing $\beta_{2}$ decreases the ratio.
+For configuration 5, increasing $a$ pushes the ratio very near 1 except for very central collisions.
 This quantity might be useful to further constrain the value of $\beta_{2}$ given that curves with small values of this parameter are well separated from those with higher values for central collisions.
 Finally in Fig.
-12, $\left(\langle(\delta(E/S))^{3}\rangle/\langle E/S\rangle^{3}\right)_{n}/\left(\langle(\delta(E/S))^{3}\rangle/\langle E/S\rangle^{3}\right)_{6}$ is shown.
-It has been claimed Jia (2022b) that this quantity could also be used to constrain $\gamma$, but the effect seems very small.
+13, $\left[\langle(\delta(E/S))^{3}\rangle/\langle E/S\rangle^{3}\right]_{i}/\left[\langle(\delta(E/S))^{3}\rangle/\langle E/S\rangle^{3}\right]_{6}$ is shown.
+It has been claimed [16] that this quantity could also be used to constrain $\gamma$, but the effect seems very small.
 Though the error bars are big, some geometrical effects can be seen, for example, introducing $\beta_{3}$ increases the ratio and diminishing $\beta_{2}$ decreases it.
-In Jia (2022a), some quantities comparable to those we computed have be obtained as well but with a different estimator for $[p_{T}]$.
-As a general rule, the effect of geometry is qualitatively similar.
+In [16], some quantities comparable to those we computed have be obtained as well but with a less efficient estimator for $[p_{T}]$. As a general rule, the effect of geometry is qualitatively similar.
 Figure 12: Same as Fig. 11 for the variance.
@@
 IV Conclusion
-We investigated the effects of a prethermalization phase in isobar collision systems (namely ${}^{96}_{44}$u +${}^{96}_{44}$u and ${}^{96}_{40}$r +${}^{96}_{40}$r ), using free-streaming as an extreme case.
+We investigated the effects of a pre-thermalization stage in isobar collision systems (namely ${}^{96}_{44}$u +${}^{96}_{44}$u and ${}^{96}_{40}$r +${}^{96}_{40}$r ), using free-streaming as an extreme case.
 The main effects anticipated were observed: the smearing out of small-scale features and reduction of spatial anisotropies (which get transferred to momentum distribution anisotropy).
@@
 Ratios of higher order eccentricities, however, show differences that can be significant when comparing with precision measurements such as the ones recorded for isobar collisions.
-Within our statistics, not much can be learned from the two-plane correlator for elliptic and triangular flows but interestingly the normalized symmetric cumulant seems sensitive to the duration of free-streaming.
+Within our statistics, not much can be learned from the two-plane correlator for elliptic and triangular flows but interestingly the normalized symmetric cumulant $\varepsilon NSC(2,3)$ seems sensitive to the duration of free-streaming.
 These studies of eccentricities yield information on the parameters $\beta_{2}$, $\beta_{3}$ and $a$.
 Furthermore, we studied the effects that nuclear geometry could have on the correlation between anisotropic flow and mean momentum to get information on $\gamma$.
-In Giacalone et al.
-(2021c), the estimator $\rho(\epsilon_{2}^{2},E/S)$ was shown to better reproduce the hydrodynamic correlator $\rho(v_{2}^{2},[p_{T}])$.
+In [52], the estimator $\rho(\varepsilon_{2}^{2},E/S)$ was shown to better reproduce the hydrodynamic correlator $\rho(v_{2}^{2},[p_{T}])$.
 Here for the first time, it is used to study isobar collisions.
-A several percent difference between the cases n=1 and 2 (with and without triaxiality) is indeed observed for $\rho_{2}$.
-In addition, and this is the main result as far as free-streaming effects are concerned, the duration of the prethermalization phase, regulated by $\tau_{FS}$, has a noticeable and diverse effect on these ratios of correlators depending on the description of the nuclear geometry: for $\rho_{2}$, it is able to enhance the separation between results with and without triaxiality effects (cases 1 and 2); for $\rho_{3}$ it generates a non-zero correlation between $\varepsilon_{3}$ and $E/S$ (proxies for $v_{3}$ and the mean transverse momentum, respectively) for the same cases.
-Another distinct effect for $\rho_{2}$ appears for case 3 (where in addition to the quadrupole deformation of cases 1 and 2, an octupole deformation is introduced), a double sign change for the largest free-streaming times ($1.35\lesssim\tau_{FS}\,({\rm fm/c})\lesssim 1.49$) appears, which can be contrasted to no or perhaps one sign change for initial condition and free-streaming time equal to $1$ fm/c.
-If verified experimentally, these features provide another tests of the predictive power of state-of-the-art hybrid hydrodynamic models Giannini et al.
-(2023).
+A several percent difference between the configurations $i=1$ and 2 (with and without triaxiality) is indeed observed for $\rho_{2}$.
+In addition, and this is the main result as far as free-streaming effects are concerned, the duration of the pre-thermalization stage, regulated by $\tau_{FS}$, has a noticeable and diverse effect on these ratios of correlators depending on the description of the nuclear geometry: for $\rho_{2}$, it is able to enhance the separation between results with and without triaxiality effects (configurations 1 and 2); for $\rho_{3}$ it generates a non-zero correlation between $\varepsilon_{3}$ and $E/S$ (proxies for $v_{3}$ and the mean transverse momentum, respectively) for the same configurations.
+Another distinct effect for $\rho_{2}$ appears for configuration 3 (where in addition to the quadrupole deformation of configurations 1 and 2, an octupole deformation is introduced), a double sign change for the largest free-streaming times ($1.35\lesssim\tau_{FS}\,({\rm fm/c})\lesssim 1.49$) appears, which can be contrasted to no or perhaps one sign change for initial condition and free-streaming time equal to $1$ fm/c.
+If verified experimentally, these features provide another tests of the predictive power of state-of-the-art hybrid hydrodynamic models [57].
 The analysis with $E/S$ as estimator for $[p_{T}]$ is also used to study the distribution of $[p_{T}]$ and its first momenta, for initial conditions only (because $E/S$ is conserved during free-streaming) for the first time.